OpenStudy (anonymous):

choose the point-slope form of the equation below that represents the line that passes through the point (-1,6) and has a slope of -3 A.y-6=-3x-3 B.y-6=-3(x+1) c.y=-3x+3 d.3x+y=3 I think it's between A-C i'm not sure about d

Nnesha (nnesha):

choose the `point slope form` what's the point slope form of a line ? :=))

OpenStudy (anonymous):

how do i get it?

Nnesha (nnesha):

oh well that's the equation `slope intercept ` form is \[\large\rm y=mx+b\]where m is slope and b is y-intercept and point slope form is \[\large\rm y-y_1=m(x-x_1)\] where m represent the slope and (x_1,y_1) the point where line passses though :))

OpenStudy (anonymous):

ok I just didn't understand so how do I get the answer.

Nnesha (nnesha):

just plug in numbers `equation that has slope of -3 and passes through the poin t( -1 ,6)` replace m with -3 and (x_1 ,y_1) point with (-1,6)

OpenStudy (anonymous):

I really don't understand it's hard for me can I get a direct answer and can you show the work so I can see how you got it It will help I will understand more like that.

Nnesha (nnesha):

there is no direct answer. which equation is **`point slope form` in the answer choices ??

Nnesha (nnesha):

point slope form is \[\large\rm y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

A

OpenStudy (anonymous):

I really don't understand

Nnesha (nnesha):

i don't know what part you didn't understand just forget about the answer this is the formula for point slope form \[\large\rm y-y_1=m(x-x_1)\] where m is slope and passes through the (x_1 ,y_1) read your question `passes through the point (-1,6) and has a slope of -3 ` slope = -3 so substitute m for -3 and passes through the (-1,6) so substitute (x_1,y_1) for (-1,6) in this order pair first represent x-coordinate so x_1 = -1 and y_1 = 6

Nnesha (nnesha):

there is only one possible answer which answer choices equation is in point slope form ??

OpenStudy (anonymous):

D?

Nnesha (nnesha):

here is an example ` line passes through the point (3,4) and has a slope of 7 ` i will replace m with 7 and x_1 with 3 and y_1 with 4 \[\rm y-4 =7(x-3)\]

OpenStudy (anonymous):

C?

Nnesha (nnesha):

don't make a guess what part you don't understand ????

Nnesha (nnesha):

why C ???

OpenStudy (anonymous):

I don't know how to plug in the numbers and solve it. I thought C because after your example it seemed more similar

Nnesha (nnesha):

this is my example \[\rm y-4 =7(x-3)\] and this is C \[y=-3x+3\] both are similar ?

Nnesha (nnesha):

as i mentioned above y=mx+b is `slope intercept ` form not point slope

Nnesha (nnesha):

plug means just replace variables with numbers just like i did in that example

OpenStudy (anonymous):

Ok so it's B? because both are showing an equation to find X and Y

Nnesha (nnesha):

formula for point slope form \[\large\rm y-y_1=m(x-x_1)\] option B \[\large y-6=-3(x+1)\]

Nnesha (nnesha):

that's the ONLY option in `point slope ` form

OpenStudy (anonymous):

Yeah thanks so much sorry for the hard time.

Nnesha (nnesha):

my pleasure :=)) just practice it's not that hard that's the easy topic in math :=))